### v2

# Minimum Bayes Risk Decoding for Error Span Detection in Reference-Free Automatic Machine Translation Evaluation

Abstract
Error Span Detection (ESD) extends automatic machine translation (MT) evaluation by localizing translation errors and labeling their severity. Current generative ESD methods typically use Maximum a Posteriori (MAP) decoding, assuming that the model-estimated probabilities are perfectly correlated with similarity to the human annotation, but we often observe higher likelihood assigned to an incorrect annotation than to the human one.
We instead apply Minimum Bayes Risk (MBR) decoding to generative ESD. We use a sentence- or span-level similarity function for MBR decoding, which selects candidate hypotheses based on their approximate similarity to the human annotation. Experimental results on the WMT24 Metrics Shared Task show that MBR decoding significantly improves span-level performance and generally matches or outperforms MAP at the system and sentence levels. To reduce the computational cost of MBR decoding, we further distill its decisions into a model decoded via greedy search, removing the inference-time latency bottleneck.This work was done during the first author’s internship at National Institute of Information and Communications Technology, Kyoto, Japan. [^1] Our models and code are available at: https://github.com/vlaks425/MBR-ESD
1 Introduction
Automatic metrics for machine translation (MT) provide a fast proxy for costly human evaluation. From early n-gram metrics such as BLEU (Papineni et al., 2002) to learned metrics such as COMET (Rei et al., 2022), correlations with human judgments have improved at the system and sentence levels. However, most metrics still output a single scalar score per sentence (Freitag et al., 2024), hiding diagnostic information such as error location and severity that is crucial for model analysis and post-editing.
Figure 1: Performance of decoding methods with Llama-3.3-70B-Inst on the WMT24 Metrics Shared Task. MBR indicates our MBR-SoftF1.
Error Span Detection (ESD) extends MT evaluation by asking metrics not only to score a translation but also to identify error spans and their severities. ESD has been part of the WMT Quality Estimation shared task since 2023 (Blain et al., 2023). Current ESD methods fall into two paradigms: token-classification approaches, such as xCOMET (Guerreiro et al., 2024), which treat ESD as per-token labeling, and generative approaches, such as GEMBA-MQM (Kocmi and Federmann, 2023), which prompt large language models (LLMs) to output MQM-style (Freitag et al., 2021a) error spans in zero- or few-shot settings.
System Error Span Annotation Log-Likelihood
Human Ich wollte fliegen, da ich ein Kind war. -10.84
Generative Model (with MBR) Ich wollte fliegen, da ich ein Kind war. -5.99
Generative Model (with MAP) Ich wollte fliegen, da ich ein Kind war. -2.86
[…]
State-of-the-art, LLM-based generative ESD methods (Freitag et al., 2024) typically rely on Maximum a Posteriori (MAP) decoding. The underlying assumption of MAP decoding is that the model-estimated probabilities are perfectly correlated with similarity to the human annotation. However, this assumption does not always hold. As illustrated in Table 1, we observe instances where the model log-likelihood is higher for an incorrect annotation produced by MAP decoding than for the correct human annotation, signifying a model error. This discrepancy, consistent with findings in other text generation tasks (Freitag et al., 2022a; Liu et al., 2022), leads us to question the optimality of MAP as a decoding objective for ESD.
Minimum Bayes Risk (MBR) decoding is a reranking method widely studied in MT (Kumar and Byrne, 2004; Tromble et al., 2008) and speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000). In this work, we apply MBR decoding to the ESD model with the objective of maximizing utility to the human annotation, independent of the model-estimated probabilities. Since the human annotation is unavailable at test time, we approximate it with a set of model-generated “support hypotheses” and select the hypothesis with the highest expected utility against this set.
We define the utility function as a similarity function over pairs of hypotheses and consider both sentence-level and span-level variants. The proposed sentence-level utility function measures similarity based on the difference between the severity-weighted error span counts of the two hypotheses. For the span-level utility function, we examine the official $\textsc{F1}(\cdot,\cdot)$ metric [^2] Referred to as the “meta-metric” in prior work. from the WMT QE Shared Task (Blain et al., 2023). However, we identify a defect in $\textsc{F1}(\cdot,\cdot)$ related to its handling of empty (error-free) annotations, which prevents it from performing soft evaluation. We therefore propose $\textsc{SoftF1}(\cdot,\cdot)$, a novel utility function that enables soft evaluation by addressing this defect.
Experimental results on the WMT24 Metrics Shared Task show that our MBR decoding, when instantiated with the proposed $\textsc{SoftF1}(\cdot,\cdot)$ utility function and applied with sufficiently large candidate sets, significantly improves span-level performance and generally matches or outperforms the MAP decoding baseline at the system and sentence levels; partial results are shown in Figure 1. To avoid the high computational cost of MBR decoding, we further distill its decisions into a model decoded via greedy search, removing the inference-time latency bottleneck.
2 Related Work
To our knowledge, our work is the first to apply reranking techniques to MT metrics.
2.1 MT Automatic Metrics
Early MT metrics such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and chrF (Popović, 2015) rely on surface-level matching and output a single scalar score per sentence. Neural metrics, including BertScore (Zhang et al., 2020), BLEURT (Sellam et al., 2020), COMET (Rei et al., 2020), and MetricX (Juraska et al., 2024), show higher correlations with human judgments (Freitag et al., 2022b; Freitag et al., 2023; Freitag et al., 2024) but usually still produce sentence-level scores only, limiting their usefulness for applications such as post-editing.
As human evaluation has shifted from scalar scores (e.g., Direct Assessment) to multidimensional annotations such as MQM (Freitag et al., 2021a), some metrics now target finer-grained evaluation, including ESD. xCOMET (Guerreiro et al., 2024) predicts error spans via token classification, while GEMBA-MQM (Kocmi and Federmann, 2023) and AutoMQM (Fernandes et al., 2023) use generative models. Generative metrics have attracted attention for their flexibility—often requiring only prompt changes—and for their competitive performance (Lu et al., 2025; Huang et al., 2024; Xu et al., 2023; Baswani et al., 2023; Feng et al., 2025; Freitag et al., 2024).
2.2 Evaluation Metrics for ESD
The alignment between model-predicted ESD annotations and the human annotation can be evaluated at the system, sentence, and span levels.
Soft Pairwise Accuracy (SPA) (Thompson et al., 2024) and Pairwise Accuracy with Tie Calibration (${\text{Acc}}_{\text{eq}}^{*}$) (Deutsch et al., 2023) measure ranking consistency between model-predicted ESD annotations and the human annotation at the system and sentence levels, respectively.
At the span-level, to our knowledge, the only existing metric for ESD is the official metric of the WMT QE Shared Task (Blain et al., 2023), $\textsc{F1}(\cdot,\cdot)$, which computes character-level precision and recall between a hypothesis and a reference annotation. This metric can also be used as an MBR utility function when the reference is a support hypothesis.
However, we identify a critical defect in $\textsc{F1}(\cdot,\cdot)$, namely its handling of empty (error-free) annotations. As detailed in Appendix A.1, when the reference annotation is empty, $\textsc{F1}(\cdot,\cdot)$ returns its minimum value for any non-empty hypothesis, regardless of how minor the error might be. This is problematic for two reasons. First, as an evaluation metric, it fails to provide a soft penalty when the human annotation is error-free. Second, as an MBR utility function, it assigns minimum utility to any non-empty candidate whenever the support hypothesis is empty. Given the diversity of support hypotheses (Kamigaito et al., 2025) and the growing prevalence of perfect translations (Freitag et al., 2024), cases with empty human annotations or empty support hypotheses are likely to be frequent. We therefore regard this behavior as a significant limitation, which motivates our soft alternative, $\textsc{SoftF1}(\cdot,\cdot)$, introduced in Section 5.2.3.
2.3 MBR Decoding
MBR decoding has a long history in speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000), word alignment (Kumar and Byrne, 2002), and statistical MT (Kumar and Byrne, 2004; Tromble et al., 2008). Recently, several studies have revisited MBR decoding for text generation, demonstrating promising results (Wu et al., 2025; Heineman et al., 2024; Lyu et al., 2025). These studies suggest that MBR decoding can help overcome some of the limitations inherent to MAP. Furthermore, the properties (Müller and Sennrich, 2021; Kamigaito et al., 2025) and efficiency (Cheng and Vlachos, 2023; Vamvas and Sennrich, 2024; Deguchi et al., 2024) of MBR decoding remain active areas of research.
3 MAP and MBR Decoding
In this section, we describe MAP and MBR decoding in the context of general conditional generative models.
Given an input $x$, a conditional generative model $M(\cdot|x)$ defines a probability distribution over the hypothesis space. Decoding for this model can be viewed as consisting of two phases: hypothesis generation and decision. In the hypothesis generation phase, a procedure such as sampling or beam search is used to produce $N$ candidate hypotheses $\mathcal{C}=\{h_{0},\ldots,h_{N-1}\}$ from the model $M(\cdot\mid x)$. In the decision phase, a decision rule assigns a score ${score}_{h}$ to each $h\in\mathcal{C}$ and returns $\mathop{\mathrm{argmax}}\limits_{h\in\mathcal{C}}score_{h}$ as the final output.
3.1 MAP Decoding
Formally, let $\mathcal{H}$ denote the entire hypothesis space. MAP decoding selects the hypothesis with the highest model-estimated probability from $\mathcal{H}$:
$h^{\mathit{MAP}}$ $=\mathop{\mathrm{argmax}}\limits_{h\in\mathcal{H}}score^{\mathit{MAP}}_{h},$
$score^{\mathit{MAP}}_{h}$ $=M(h\mid x).$
Finding the global maximizer over $\mathcal{H}$ is typically intractable due to the enormous size of the space. In practice, MAP decoding is approximated by restricting the search to a finite set of candidate hypotheses $\mathcal{C}\subset\mathcal{H}$:
$h^{\mathrm{MAP}}\approx\mathop{\mathrm{argmax}}\limits_{h\in\mathcal{C}}M(h\mid x).$
3.2 MBR Decoding
Unlike MAP, MBR explicitly optimizes task utility rather than model-estimated probability. Let $\mathcal{Y}$ denote the space of (unknown) ground-truth output $y$ for input $x$ and let $u(\cdot,\cdot)$ be a utility function, MBR decoding selects the hypothesis that minimizes the expected risk—or equivalently, maximizes the expected utility—with respect to $\mathcal{Y}$:
$h^{\mathit{MBR}}$ $=\mathop{\mathrm{argmax}}\limits_{h_{c}\in\mathcal{C}}score^{\mathit{MBR}}_{h},$
$score^{\mathit{MBR}}_{h}$ $={\mathbb{E}_{y\in\mathcal{Y}}\left[u(h,y)\mid x\right]}.$ (1)
Since $y$ is unknown in practice, MBR decoding approximates this expectation using a set of “support hypotheses” $\mathcal{S}\subseteq\mathcal{H}$ drawn from $M(\cdot|x)$:
$score^{\mathit{MBR}}_{h}$ $\approx{\mathbb{E}_{h_{s}\in\mathcal{S}}\left[u(h,h_{s})\mid x\right]},$ (2)
$\approx\sum_{h_{s}\in\mathcal{S}}u(h,h_{s}){M}(h_{s}\mid x).$
A common practice is to use the same set for candidate and support hypotheses ($\mathcal{C}=\mathcal{S}$) and to assume a uniform distribution over support hypotheses, yielding a simple average utility:
$score^{\mathit{MBR}}_{h}$ $\approx\frac{1}{|\mathcal{S}|}\sum_{h_{s}\in\mathcal{S}}u(h_{c},h_{s}).$ (3)
4 Error Span Detection with MAP Decoding
4.1 Generative ESD Task
We focus on the ESD task in the reference-free setting: given an input $x=(s,t)$ consisting of a source sentence $s$ and its translation $t$, the model outputs an error annotation (a set of error spans with associated severities).
Let $E$ denote a hypothesis containing a finite number of major and minor error spans, and let $\mathcal{H}_{\text{ESD}}$ represent the space of all admissible hypotheses. We denote the distribution of the (unknown) ground-truth human annotation as $\mathcal{Y}_{\text{ESD}}$. The generative ESD model, $M_{\text{ESD}}(E|x)$, estimates the probability of a hypothesis $E\in\mathcal{H}_{\text{ESD}}$.
For notation, we use $E_{c}$ to denote a candidate hypothesis, $E_{s}$ for a support hypothesis, and $E_{y}$ for the human annotation. Similarly, $\mathcal{C}_{\text{ESD}}$ and $\mathcal{S}_{\text{ESD}}$ represent the sets of candidate and support hypotheses, respectively.
4.2 MAP Decoding for ESD
Under MAP decoding, the decision score for a hypothesis $E\in\mathcal{C}_{\text{ESD}}$ is defined as its model-estimated probability:
$score^{\mathit{MAP}}_{E}=M_{\text{ESD}}(E\mid x).$
As discussed in §1, the MAP objective implicitly assumes that model-estimated probabilities are perfectly correlated with similarity to the human annotation. However, as illustrated by the failure case in Table 1, this assumption frequently does not hold for ESD. This discrepancy motivates the adoption of MBR decoding as a robust alternative.
5 Proposed Method
5.1 MBR Decoding for ESD
Figure 2: An overview of our MBR decoding for generative ESD models. Error spans are highlighted in red.
We propose applying MBR decoding to generative ESD models to address the limitations of MAP decoding. Figure 2 provides an overview of our method.
Analogous to Eq. 1, the MBR score for a hypothesis $E$ is defined as:
$score^{\mathit{MBR}}_{E}$ $={\mathbb{E}_{E_{y}\in\mathcal{Y}_{\text{ESD}}}\left[u(E,E_{y})\mid x\right]}.$ (4)
As in Eq. 3, we approximate this expectation using a finite sample $\mathcal{S}_{\mathrm{ESD}}\subset\mathcal{H}_{\mathrm{ESD}}$:
$score^{\mathit{MBR}}_{E}$ $\approx\frac{1}{|\mathcal{S}_{\text{ESD}}|}\sum_{E_{s}\in\mathcal{S}_{\text{ESD}}}u(E,E_{s}).$ (5)
We specify the utility functions $u(\cdot,\cdot)$ in the next section.
5.2 Utility Functions of MBR Decoding in ESD
In this section, we describe three instantiations of MBR decoding for generative ESD models. The utility functions fall into two categories: sentence and span level.
Notably, although SPA and ${\text{Acc}}_{\text{eq}}^{*}$ (mentioned in §2.2) are standard metrics for ESD, they cannot serve as MBR utility functions. This is because they measure ranking consistency over a dataset rather than computing a similarity score for a single pair of error annotations.
5.2.1 MBR-ScoreSim
We introduce MBR-ScoreSim, which employs our proposed $\textsc{ScoreSim}(\cdot,\cdot)$ as the utility function. $\textsc{ScoreSim}(\cdot,\cdot)$ is a sentence-level utility function that measures the similarity between two hypotheses based on a severity-weighted error span count.
Let $N^{\text{maj}}(\cdot)$ and $N^{\text{min}}(\cdot)$ denote the counts of major and minor error spans in $E$, respectively. Following common practice (Freitag et al., 2021a; Juraska et al., 2024), we first convert the error spans into a scalar sentence-level score using the standard MQM scoring function:
$\textsc{Score}(E)=\max(w_{\text{maj}}N^{\text{maj}}(E)+w_{\text{min}}N^{\text{min}}(E),\alpha).$
Here $w_{\text{maj}}<0$ and $w_{\text{min}}<0$ are severity-dependent penalties, and $\alpha<0$ is a lower bound on the total penalty.
We then define the similarity between $E_{c}$ and $E_{s}$ as one minus the normalized absolute difference between their scores, using $|\alpha|$ for normalization to yield a value in $[0,1]$:
$\textsc{ScoreSim}(E_{c},E_{s})=1-\frac{|\,\textsc{Score}(E_{c})-\textsc{Score}(E_{s})\,|}{|\alpha|}.$
5.2.2 MBR-F1
For span-level utility, we examine MBR-F1, which employs the standard $\textsc{F1}(\cdot,\cdot)$ metric from the WMT QE Shared Task. This function calculates the harmonic mean of character-level precision and recall between two annotations (see Appendix A.1 for details).
While $\textsc{F1}(\cdot,\cdot)$ is widely used, it exhibits a defect when handling empty annotations. Specifically, if the support hypothesis $E_{s}$ is empty (i.e., contains no errors), $\textsc{F1}(\cdot,E_{s})$ assigns the minimum utility (0) to any non-empty candidate $E_{c}$, regardless of how minor the error in $E_{c}$ might be (see Appendix A.1 for the derivation). This “all-or-nothing” behavior fails to capture the nuance that a hypothesis with a single minor error is arguably better—and more similar to an error-free translation—than one with multiple major errors.
5.2.3 MBR-SoftF1
We introduce MBR-SoftF1, an instance of MBR decoding using our proposed $\textsc{SoftF1}(\cdot,\cdot)$ as the utility function. $\textsc{SoftF1}(\cdot,\cdot)$ provides a continuous, “soft” evaluation of mismatch, ensuring robustness even when comparing against empty support hypotheses.
Formally, let $L$ be the number of characters in translation $t$. An error annotation $E$ consists of two index sets $E^{maj},E^{min}\subseteq\{1,\dots,L\}$, representing the character positions of major and minor errors. We represent this annotation not as discrete sets, but as a dense severity vector $\mathbf{v}\in\mathbb{R}^{L}$. For each character position $i\in\{1,\dots,L\}$, the vector element $\mathbf{v}^{(i)}$ represents the aggregated severity: $\mathbf{v}^{(i)}=\beta\cdot\mathbb{I}[i\in E^{maj}]+\gamma\cdot\mathbb{I}[i\in E^{min}]$, where $\beta$ and $\gamma$ represent the penalties for major and minor errors, respectively. The total error severity of an annotation is the $\ell_{1}$ norm of its vector, $\|\mathbf{v}\|_{1}=\sum_{i=1}^{L}|\mathbf{v}^{(i)}|$.
We measure the discrepancy between a candidate hypothesis $E_{c}$ (with vector $\mathbf{v}_{c}$) and a support hypothesis $E_{s}$ (with vector $\mathbf{v}_{s}$) using the $\ell_{1}$ distance, $\|\mathbf{v}_{c}-\mathbf{v}_{s}\|_{1}$. Based on these vector representations, we define Soft Precision (SoftP) and Soft Recall (SoftR) as:
$\textsc{SoftP}(E_{c},E_{s})$ $=1-\frac{\|\mathbf{v}_{c}-\mathbf{v}_{s}\|_{1}}{L+\|\mathbf{v}_{c}\|_{1}},$
$\textsc{SoftR}(E_{c},E_{s})$ $=1-\frac{\|\mathbf{v}_{c}-\mathbf{v}_{s}\|_{1}}{L+\|\mathbf{v}_{s}\|_{1}}.$
[…]
Unlike the standard $\textsc{F1}(\cdot,\cdot)$, $\textsc{SoftF1}(\cdot,\cdot)$ returns a continuous value reflecting the degree of mismatch. Crucially, even if the support hypothesis is empty ($\mathbf{v}_{s}=\mathbf{0}$), a candidate with a small error vector $\mathbf{v}_{c}$ will yield a high utility score closer to $1$, rather than collapsing to 0 (see proof in Appendix A.2). This property enables MBR decoding to meaningfully rank candidates even when the consensus among support hypotheses is that the translation is error-free. Furthermore, $\textsc{SoftF1}(\cdot,E_{y})$ can also serve as an effective evaluation metric for ESD.
6 Experiments on MBR Decoding
Llama-3.3-70B-Inst Gemma2-27b-it
NN Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1 SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
1 Greedy Search .820 .567 .922 .475 .816 .524 .900 .415
16 MAP .831 .565 .921 .537 .806 .529 .934 .571
MBR-ScoreSim   .855   .573 .906 .381 .802   .552 .890 .252
MBR-F1 .839 .566 .918 .522 .810 .538 .924 .561
MBR-SoftF1 .845 .565 .929 .510   .832 .531 .935 .545
64 MAP .831 .568 .922 .539 .797 .521 .933 .568
MBR-ScoreSim   .849   .578 .902 .347 .814   .557 .880 .197
MBR-F1 .841 .568 .913 .529 .821   .537 .922   .575
MBR-SoftF1 .843 .569   .931 .514   .833   .533   .938 .554
256 MAP .823 .568 .919 .531 .844 .519 .926 .545
MBR-ScoreSim .843   .579 .898 .328 .815   .560 .873 .170
MBR-F1   .844 .569 .923 .532 .814   .539 .916   .579
MBR-SoftF1   .848   .571  .932 .513 .846   .532   .938 .553
Table 2: Evaluation results for the WMT 2024 Metrics Shared Task, conducted with Llama-3.3-70B-Inst and Gemma2-27b-it. We employ SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1 as metrics at the system, sentence, and span-levels, respectively. We report averaged scores across all translation directions. The best for the same model with NN is in bold. indicates significantly better performance than the MAP baseline with the same NN (# of candidate hypotheses) on all translation directions ($p<0.05$).
6.1 Experimental Setup
Datasets and Models
[…]
Hypothesis Generation and Decision
For hypothesis generation, the model generates $N\in\{16,64,256\}$ hypotheses for each input, with Top-$K$ ($K=10$) sampling (Fan et al., 2018) and a temperature of $2.0$. This high temperature promotes diversity among candidate hypotheses, which is expected to indirectly elevate utility variance and thereby reduce utility estimation error (Kamigaito et al., 2025).
For the decision phase, we compared the MAP baseline against the three MBR decoding methods introduced in §5.2. Following prior work (Freitag et al., 2021a; Juraska et al., 2024; Blain et al., 2023), we merged “critical” errors into the “major” category and adopt their recommended parameter values for Eq. 6: $w_{\text{maj}}=-5$, $w_{\text{min}}=-1$, $\alpha=-25$, $\beta=1$ and $\gamma=0.5$.
Evaluation
Following the WMT Metrics Shared Task, we used SPA (Thompson et al., 2024) and ${\text{Acc}}_{\text{eq}}^{*}$ (Deutsch et al., 2023) for system- and sentence-level evaluation, respectively. For the span level, we used $\textsc{SoftF1}(\cdot,E_{y})$ as the metric and reported the average scores over test instances. We selected $\textsc{SoftF1}(\cdot,E_{y})$ over $\textsc{F1}(\cdot,E_{y})$ as we consider it a more robust metric for ESD, particularly for handling empty annotations. For the SPA and ${\text{Acc}}_{\text{eq}}^{*}$, we employed the PERM-BOTH test (Deutsch et al., 2021). For SoftF1, [^4] SoftF1 indicates the average of the return values by $\textsc{SoftF1}(\cdot,E_{y})$ during evaluation across the entire dataset. Similar definitions apply to other functions. , since it operates at the instance level, we used paired bootstrap resampling (Koehn, 2004). We used MAP as the baseline for all statistical significance tests. For all ESD methods, sentence-level scores were derived using Eq. 6, system-level scores were calculated by averaging all sentence-level scores within that MT system.
6.2 Experimental Results
Phi-4 Mistral-Large-Inst
NN Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1 SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
1 Greedy Search .825 .548 .792 .307 .815 .538 .774 .171
16 MAP .817 .553 .920 .533 .819 .541 .913 .463
MBR-ScoreSim   .837 .554 .893 .284 .825 .530 .865 .119
MBR-F1 .822 .528 .925   .563   .834 .536 .885 .454
MBR-SoftF1   .847 .535   .936   .552 .828 .524   .930   .484
64 MAP .822 .551 .922 .550 .825 .530 .914 .471
MBR-ScoreSim   .852   .559 .881 .238 .832   .549 .857 .107
MBR-F1 .831 .530 .924   .575 .830 .536 .879   .498
MBR-SoftF1   .848 .535   .938 .559 .828 .528   .933   .496
256 MAP .832 .545 .921 .549 .841 .525 .905 .427
MBR-ScoreSim .843   .563 .875 .222 .831   .558 .852 .104
MBR-F1 .841 .536 .918   .577 .834   .538 .819   .506
MBR-SoftF1   .850 .549   .938   .558 .846 .528   .934   .494
Table 3: Evaluation results for the WMT 2024 Metrics Shared Task, conducted with Phi-4 and Mistral-Large-Inst. The element definitions in this table are the same as those in Table 2.
Drawing on the results presented in Table 2 and 3, we compare the performance of our proposed MBR instantiations against the MAP baseline. Our key findings are detailed below.
Performance of MBR-ScoreSim
At $N\geq 64$, MBR-ScoreSim achieved significantly higher ${\text{Acc}}_{\text{eq}}^{*}$ scores compared to MAP, but lacks consistent improvement on SPA and SoftF1. This indicates a trade-off: MBR-ScoreSim is highly effective at the specific level it optimizes (sentence-level) but struggles to generalize to system or span-level improvements. Nevertheless, it remains valuable for applications that rely on sentence-level metrics, such as reranking for MT.
Performance of MBR-F1
MBR-F1 performed comparably to MAP with only marginal fluctuations. We attribute this stagnation to the empty-annotation defect in F1(,)($\cdot$$\cdot$)5.2.2), which hinders its effectiveness as an optimization objective and confirms the need for our soft alternative.
Performance of MBR-SoftF1
At $N=256$, MBR-SoftF1 consistently improves SoftF1 over MAP for all four models, with these improvements being statistically significant in every setup; for SPA and ${\text{Acc}}_{\text{eq}}^{*}$, MBR-SoftF1 typically matches or outperforms MAP, although these gains are not always statistically significant. For example, with Mistral-Large-Inst, the SPA and ${\text{Acc}}_{\text{eq}}^{*}$ scores under MBR-SoftF1 are numerically close to those of MAP and the differences are not significant. At $N=16$, MBR-SoftF1 performed comparably to MAP with only marginal fluctuations, highlighting a key limitation: MBR-SoftF1 effectively trades computational cost for quality, requiring large sets of hypotheses ($N$). To address this resource dependency, we investigate MBR distillation in §8.
Conclusion
While MBR-ScoreSim and MBR-F1 are constrained by optimization trade-offs or defects in their utility functions, MBR-SoftF1 largely mitigates these issues when $N$ is sufficiently large. Overall, these results support the conclusion that MBR decoding, instantiated with MBR-SoftF1 and a sufficiently large candidate set, significantly improves span-level performance and generally matches or outperforms MAP at system and sentence levels. Finally, we note a limitation regarding span-level conclusions, which depend on $\textsc{SoftF1}(\cdot,\cdot)$ being a reliable span-level metric. Since MBR-SoftF1 was directly optimized by $\textsc{SoftF1}(\cdot,\cdot)$ as the utility function, the evaluation metric intrinsically favors this method, a phenomenon known as metric bias (Kovacs et al., 2024). Lacking a neutral metric, we leave further investigation of this bias to future work.
7 Analysis of MBR Decoding
7.1 MBR Performance Upper Bound
Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
Model-ScoreSim .843 .579 .898 .328
Model-F1 .844 .569 .923 .532
Model-SoftF1 .848 .571 .932 .513
Oracle-ScoreSim .925 .926 .943 .701
Oracle-F1 .906 .776 .970 .843
Oracle-SoftF1 .902 .845 .981 .811
Table 4: Understanding the performance upper bound of MBR through the oracle setup. Analysis is conducted under the sub-setup of Llama-3.3-70B-Inst with $N=256$. Model-* indicates MBR based on model-generated support hypotheses, while Oracle-* indicates MBR based on the human annotation.
To understand the upper bound on MBR performance, we conducted an oracle experiment. We evaluated the discrepancy between two MBR decoding strategies: (1) Oracle MBR, which computes utility against the human annotation (Eq. 4) [^5] We used the human annotation from the WMT Metrics Shared Task dataset, provisionally assuming that the probability of the human annotation for each $x$ is $1$., and (2) Standard MBR, which relies on model-generated support hypotheses (Eq. 5).
As shown in Table 4, Oracle MBR significantly outperforms the standard MBR (based on support hypotheses) across all metrics. This substantial gap indicates that the quality of the support hypotheses is a limiting factor, suggesting significant room for improvement in future work.
7.2 MBR-SoftF1 Scalability
MBR decoding effectively leverages increased test-time computation to improve performance, a paradigm often referred to as test-time scaling (Snell et al., 2024). Consequently, we investigate its scalability with respect to $N$.
NN Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
256 MAP .823 .568 .919 .531
MBR-SoftF1 .848 .571 .932 .513
1024 MAP .831 .567 .916 .524
MBR-SoftF1 .849 .571 .934 .513
Table 5: Analysis of MBR-SoftF1 scalability. Analysis is conducted under the sub-setup of Llama-3.3-70B-Inst.
Due to computational constraints, we performed this analysis exclusively on the Llama-3.3-70B-Inst sub-setup, increasing $N$ to $1024$. As presented in Table 5, MBR-SoftF1 shows negligible performance gains at $N=1024$ compared to $N=256$ on SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1. We hypothesize that the model fails to generate sufficient diversity among candidates with the current temperature and Top-$K$ settings to benefit from the larger sample size. While tuning these hyperparameters for larger $N$ could potentially unlock further gains, we leave this investigation for future work.
7.3 Number and Proportion of Error Spans
Method Major Minor $\frac{\text{Major}}{\text{Minor}}$
Human 9.0K 15.5K 0.58
MAP 18.9K 18.7K 1.01
MBR-ScoreSim 13.0K 59.6K 0.21
MBR-F1 20.5K 25.9K 0.79
MBR-SoftF1 13.1K 20.1K 0.65
Table 6: Analysis of the error span distribution across different ESD annotators. Generative ESD metrics are based on Llama-3.3-70B-Inst with $N=256$.
We analyzed the distribution of error spans in the outputs of human annotators versus generative ESD metrics. The results are summarized in Table 6. We observe that MAP tends to over-generate error spans compared to human annotators. Furthermore, it exhibits a bias in severity proportion: $\frac{\text{Major}}{\text{Minor}}$ for MAP ($1.01$) deviates significantly from the human reference ($0.58$).
Notably, although the utility functions for MBR-F1 and MBR-SoftF1 do not explicitly optimize for error counts, they effectively mitigate this proportional bias, bringing $\frac{\text{Major}}{\text{Minor}}$ much closer to human levels. Moreover, MBR-SoftF1 produces absolute counts of major and minor errors that are closer to the human reference. Since the impact of this discrepancy in error distribution on downstream tasks remains an open question, we did not explore direct interventions in this study. However, a potential remedy for future work would be to incorporate a term related to error span counts into the utility function to explicitly control the distribution.
7.4 Comparison with State-of-the-Art
Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
Llama-MAP .823 .568 .919 .531
Llama-MBR-SoftF1 .848 .571 .932 .513
xCOMET-ESD .757 .553 .889 .302
xCOMET-Reg .844 .581 - -
xCOMET-QE-ESD .688 .541 .879 .289
xCOMET-QE-Reg .825 .549 - -
Table 7: Comparison of MBR-SoftF1 (with Llama-3.3-70B-Inst and $N=256$) vs. xCOMET. *-Reg and *-ESD denote regression (score-only) and token-classification modes, respectively. xCOMET-QE-* indicates a reference-free variant; others are reference-based.
Table 7 shows the comparison between generative metrics with MBR-SoftF1 and the state-of-the-art ESD metric (Freitag et al., 2023; Freitag et al., 2024), xCOMET (Guerreiro et al., 2024) (Unbabel/XCOMET-XXL). According to SPA and SoftF1, Llama-MBR-SOFTF1 (i.e. MBR-SoftF1 with Llama-3.3-70B-Inst) outperforms any variants of xCOMET, ranking lower than xCOMET-Reg only on ${\text{Acc}}_{\text{eq}}^{*}$. While these results suggest that the generative ESD method with MBR decoding is competitive for the ESD task, it is worth noting that it entails much higher computational overhead than xCOMET.
8 Experiment on MBR Distillation
Despite the significant performance gains demonstrated in §6, MBR decoding comes at a high computational cost. Specifically, it relies on generating and scoring a large set of hypotheses (i.e., a large $N$), which incurs substantial latency and compute overhead during inference. To alleviate this bottleneck, we investigate MBR distillation, a technique designed to transfer the superior performance of MBR decoding into the model’s weights, thereby enabling comparable results using efficient greedy search.
Following prior work on MBR distillation (Finkelstein and Freitag, 2024; Wang et al., 2025; Yang et al., 2024), we employ Direct Preference Optimization (DPO) (Rafailov et al., 2023) to train the model to approximate MBR outputs. DPO is a pairwise ranking optimization method typically used to align models with human preferences by maximizing the margin between a preferred and a rejected output. We adapt DPO for MBR distillation by constructing a pairwise dataset where the candidate hypothesis with the highest utility is treated as the preferred output, and the one with the lowest utility serves as the rejected output.
8.1 Experimental Setup
Datasets and Model
The training dataset was derived from $(s,t)$ pairs in the WMT20–22 Metrics Shared Task datasets (Mathur et al., 2020; Freitag et al., 2021b; Freitag et al., 2022b). We used the WMT24 Metrics Shared Task as the test set. We used Llama-3.3-70B-Inst as the model. To generate the training targets, we employed the same MBR setup as in §6: $N=256$ with MBR-SoftF1. We refer to the model distilled via DPO as Distill-Greedy, which employs standard greedy decoding at inference time. Model training details are provided in Appendix C.
Evaluation
The evaluation metrics for ESD are identical to those used in §6.
8.2 Experimental Results
N Method SPA ${\text{Acc}}_{\text{eq}}^{*}$ SoftF1 F1
1 Greedy Search .820 .567 .922 .475
256 MAP .823 .568 .919 .531
MBR-SoftF1 .848 .571 .932 .513
1 Distill-Greedy .857 .572 .938 .538
Table 8: Comparison of MBR-SoftF1 (with $N=256$) and the distilled model (Distill-Greedy, with greedy search). Results other than Distill-Greedy are cited from Table 2. Experiments are conducted using Llama-3.3-70B-Inst on the WMT24 Metrics Shared Task. Distill-Greedy denotes the model distilled from MBR-SoftF1 outputs via DPO.
As evidenced in Table 8, Distill-Greedy achieves scores nearly identical to MBR-SoftF1 on SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1. This result confirms that our distilled model using greedy search can match the performance of full MBR decoding while eliminating the dependency on large $N$. Given that the computational cost of the distilled model with greedy search is substantially lower than that of the pre-distillation model with MBR decoding, we conclude that MBR distillation offers a highly promising pathway toward realizing efficient, high-performance generative ESD methods.
9 Conclusions and Future Work
In this work, we introduced MBR decoding for generative ESD models, marking the first application of reranking techniques to automatic MT metrics. Extensive experimental results show that our MBR decoding, when instantiated with MBR-SoftF1 and applied with sufficiently large candidate sets, significantly improves span-level performance and generally matches or outperforms MAP at system and sentence levels. Furthermore, we established that MBR distillation allows a model decoded via greedy search to match the performance of MBR, effectively resolving the inference latency bottleneck.
We believe the following are promising directions for future work: (1) Achieving further performance gains by integrating multiple utility functions. (2) Applying MBR decoding to self-improvement training in the ESD model.
Appendix A $\textsc{F1}(\cdot,\cdot)$ vs. $\textsc{SoftF1}(\cdot,\cdot)$
A.1 $\textsc{F1}(\cdot,\cdot)$
For each character position $i\in[1,L]$, we define binary indicators for the presence of major or minor errors: $m^{i}=\mathbb{I}[i\in E^{maj}],\qquad n^{i}=\mathbb{I}[i\in E^{min}]$, and these indicators for the hypothesis and reference as $(m_{c}^{i},n_{c}^{i})$ and $(m_{s}^{i},n_{s}^{i})$, respectively. The character-level true-positive credit, $c^{i}\in\{0,\gamma,\beta\}$, is calculated based on the agreement between severities. Assuming $\beta$ is the weight for major errors and $\gamma$ for minor errors (typically $\beta>\gamma$): $c^{i}=\max(\beta m_{c}^{i}m_{s}^{i},\beta n_{c}^{i}n_{s}^{i},\gamma m_{c}^{i}n_{s}^{i},\gamma n_{c}^{i}m_{s}^{i}).$ We denote the total severity score of an annotation $E$ as $S(E)$: $S(E)=\sum_{i=1}^{L}\max(m^{i},n^{i})$, then precision P and recall R are defined as:
$\textsc{P}=\begin{cases}1,&\text{if }S(E_{c})=S(E_{s})=0,\\ \frac{\sum_{i}c^{i}}{S(E_{c})},&\text{if }S(E_{c})>0,\\ 0,&\text{otherwise,}\end{cases}\textsc{R}=\begin{cases}1,&\text{if }S(E_{c})=S(E_{s})=0,\\ \frac{\sum_{i}c^{i}}{S(E_{s})},&\text{if }S(E_{s})>0,\\ 0,&\text{otherwise.}\end{cases}$
[…]
A Defect in $\textsc{F1}(\cdot,\cdot)$
If $E_{s}=\emptyset$, then $m_{s}^{i}=n_{s}^{i}=0$, and hence $c^{i}=\max(\beta m_{c}^{i}0,\beta n_{c}^{i}0,\gamma m_{c}^{i}0,\gamma n_{c}^{i}0)=0$, for all $i$, so $\sum_{i}c^{i}=0$. Thus, when $E_{c}=\emptyset$, it perfectly matches $E_{s}$, and $\textsc{F1}(\cdot,\cdot)$ correctly returns $1$. However, if $E_{c}\neq\emptyset$, even if it marks only a single character as a minor error (i.e. very close to empty), the score collapses to 0.
A.2 $\textsc{SoftF1}(\cdot,\cdot)$: Non-empty vs. Empty Spans
If $E_{c}\neq\emptyset$ and $E_{s}=\emptyset$, then:
$\textsc{SoftP}(E_{c},E_{s})=1-\frac{\|\mathbf{v}_{c}\|_{1}}{L+\|\mathbf{v}_{c}\|_{1}}=\frac{1+L}{L+\|\mathbf{v}_{c}\|_{1}}>0,~~~\textsc{SoftR}(E_{c},E_{s})=1-\frac{\|\mathbf{v}_{c}\|_{1}}{L}>0.$
which implies $\textsc{SoftF1}(E_{c},E_{s}=\emptyset)>0$. This behavior differs from $\textsc{F1}(E_{c},E_{s}=\emptyset)$ and highlights a desirable property of $\textsc{SoftF1}(E_{c},E_{s})$: when only one of $E_{c}$ and $E_{s}$ is empty, it does not simply return 0 or $1$, but a decimal value between 0 and $1$ that reflects the degree of the match.
Appendix B Details of Hypothesis Generation and Decision
We used guided generation (Willard and Louf, 2023) to ensure that each hypothesis adheres to the specified JSON format.
[…]
Appendix C Training Details of MBR Distillation
We used the Lion optimizer (Chen et al., 2023) with a learning rate of $10^{-7}$, a warmup step of $50$, a weight decay of $0.1$, a batch size of $256$, and train for a maximum of $5$ epochs. We set the regularization term weight of DPO to $0.5$. We trained the model in BF16 precision on a machine equipped with 8 NVIDIA H200 GPUs. We used DeepSpeed (Rajbhandari et al., 2020) Zero3 for multi-GPU training. Our training codebase is based on the TRL library (von Werra et al., 2020).
The dataset comprises a total of $110,699$ samples. We randomly split the data, reserving $10\%$ for the validation set and using the remainder for the training set. Upon completion of training, the checkpoint with the lowest loss on the validation set was selected as the final model.
[…]